# Rapid, strongly magnetized accretion in the zero-net-vertical-flux shearing box

### v1

@@ Appendix C The effect of the density floor
Most simulations in the main text impose a floor on the density at $\rho_{\rm flr}=10^{-4}$ of the initial central density, which is important for stabilizing the numerical scheme and allowing somewhat larger timesteps. While undesirable, this is needed because the large $v_{\rm A}$ and highly supersonic turbulence at large $z$ cause frequent code crashes unless a very conservative CFL number is used (${\lesssim}0.1$), significantly increasing the cost of the simulations (which are already much more expensive than those in the high-$\beta$ regime). While the low-$\beta$ simulations studied in the main text do not show any obvious effects and mostly have a minimum mean density remains a factor of $\simeq 2$-$3$ above the chosen floor, it is important to check that the maintenance of the low-$\beta$ state does not depend on $\rho_{\rm flr}$.

### v3

@@ Appendix C The effect of the density floor
Most simulations in the main text impose a floor on the density at $\rho_{\rm flr}=10^{-4}$ of the initial central density, which is important for stabilizing the numerical scheme and allowing somewhat larger timesteps. While undesirable, this is needed because the large $v_{\rm A}$ and highly supersonic turbulence at large $z$ cause frequent code crashes unless a very conservative CFL number is used (${\lesssim}0.1$), significantly increasing the cost of the simulations (which are already far more expensive than those in the high-$\beta$ regime). While the low-$\beta$ simulations studied in the main text do not show any obvious effects and mostly have a minimum mean density that remains a factor of $\simeq 2$-$3$ above the chosen floor, it is important to check that the maintenance of the low-$\beta$ state does not depend on $\rho_{\rm flr}$.
@@
Figure 20: Space-time butterfly plot showing $\overline{B_{y}}$ for simulation $\beta$R0.1-NVF2-norot, which has a $\beta_{z}=100$ vertical flux, but with rotational effects removed. The system does not sustain the strong $\beta\lesssim 1$ azimuthal field that occurs with rotation (cf. Fig. 9).
Appendix D The necessity of shear and rotation
To help diagnose the mechanism(s) by which the low-$\beta$ state is maintained, in Fig. 19 we rerun two cases that transition into the low-$\beta$ state in the standard shearing box ($\beta$R0.1 and $\beta$0.01), but selectively remove either shear or rotation. The purpose of these modifications is not to explore physically relevant scenarios, but rather as numerical experiments to probe the important physics. Removing the shear simply involves setting $q=0$ in Eqs. 1a to 1c, while removing rotation is achieved by dropping the first two terms on the second line of Eq. 1b, or equivalently, by taking the limit $\Omega\rightarrow 0$ with $q\Omega=S={\rm const.}$ Both cases retain the vertical stratification via $-\rho\Omega^{2}z\hat{\bm{z}}$.
The top two panels of Fig. 19 show the effect of removing the mean shear flow; these clearly do not maintain interesting dynamics. The flux continuously escapes in $\beta$R0.1-noshear, while $P_{B}$ becomes flat in $\beta$0.01 due to its Outflow boundary conditions, and both develop perfect thermal profiles in the midplane. The detailed structure at larger $z$ is affected by the density floor and so is untrustworthy, but the obvious difference with Fig. 2 nonetheless clearly demonstrates the importance of the shear.
As disussed in §4.1, removing rotation kills the MRI, as well as changing the turbulence properties and Parker instability. The lower panels of Fig. 19 show $\beta$R0.1-norot and $\beta$0.01-norot in which the rotation is artificially removed. Neither of these maintain the low-$\beta$ state, developing a thermally dominated midplane. Interestingly, dynamo cycles develop in $\beta$R0.1-norot, despite the lack of the MRI. Assuming these are of similar origin to those in the high-$\beta$ state (given they have a similar period and structure) it is clear that the nomenclature ‘MRI dynamo cycles’ is not accurate. Given this system’s small $H_{\rm th}$ and the density floor, this would be worth a dedicated study in future work. In the lower-$\beta$ case ($\beta$0.01-norot) we do not see this feature, presumably due to the poorly resolved midplane and outflow boundary conditions that lead to $P_{B}\approx{\rm const.}$ at $z\gtrsim 2H_{\rm th}$. Notwithstanding, given the clear difference with $\beta$0.01, we can conclude that rotation is crucial for maintaining the low-$\beta$ state.
In Fig. 20, we remove the rotation for a case with a strong vertical field $\beta_{z}=100$. This causes the system to lose its azimuthal flux, generating cyclic behavior in $\overline{B_{y}}$ and a thermal midplane. This is similar to the high-$\beta$ ZNVF system (cf. Fig. 19) although with somewhat stronger mean-field cycles. The behavior is consistent with our interpretation in §5.4 that the NVF low-$\beta$ state is effectively a single large-scale MRI mode, which requires rotation to sustain itself. It also supports the idea that the dynamo cycles are not a consequence of the MRI, which does not exist in the absence of rotation.